# Polynomial-order oscillations in geometric discrepancy

### v1

@@ 3 The second method
The harmonic analysis arguments we are to employ in this section have been recently developed in [MR4358540] and [beretti2025fouriertransformplanarconvex]. These make use of a geometric quantity that we define as follows.
@@
Let $C\subset\mathbb{R}^{2}$ be a convex body. For an angle $\theta\in[0,2\pi)$ and value $\lambda>0$, we define the chord of $C$ in direction $\mathbf{u}(\theta)=(\cos\theta,\sin\theta)$ at distance $\lambda$ as
@@
One may notice that the decay of $|K_{C}(\theta,\lambda)|$ as $\lambda\to 0^{+}$ much depends on the local geometry of the (if just one) point $b\in\partial C$ whose inner normal is $\mathbf{u}(\theta)$. As examples, if $\partial C$ is $\mathcal{C}^{2}$ and has positive curvature at $b$, then $|K_{C}(\theta,\lambda)|$ ultimately decays as $\lambda^{1/2}$, while if $b$ is the vertex at a corner, then $|K_{C}(\theta,\lambda)|$ ultimately decays as $\lambda$.
@@
We refer to [beretti2025fouriertransformplanarconvex, Lem. 3.3] for a detailed proof of the latter, but we mention that it relies upon the one-dimensional results in [MR1166380] and [MR4358540].
@@
${\mu}=\sum_{\mathbf{p}\in\mathcal{P}}\mu_{\rm D}(\mathbf{p})-N\mu_{\rm H},$
@@
$\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+C)=\int_{\mathbb{T}^{2}}\mathfrak{P}\{\mathds{1}_{C}\}({\boldsymbol{\tau}}-\mathbf{x})\,\mathrm{d}{\mu}(\mathbf{x})=\left(\mathfrak{P}\{\mathds{1}_{C}\}\ast{\mu}\right)(\boldsymbol{\tau}).$
@@
the function of the Fourier coefficients of $f$. In particular, we notice that it holds $\mathcal{F}\{{\mu}\}(\mathbf{0})=0$, and moreover, for every $\mathbf{n}\in\mathbb{Z}^{2}$, it holds
@@ 3.1 Geometric construction and Fourier estimates
We start our construction. Consider a planar convex body $C$ such that it has a centre of symmetry (this ensures $\gamma_{C}=|K_{C}|$) and is symmetric with respect to the $y$-axis. Then, suppose that $\mathbf{0}\in\partial C$ and that its inner normal is $(0,1)$. By symmetry, we may construct $\partial C$ in the first quadrant $\mathbb{R}_{+}^{2}$. To this purpose, consider the graphs in $\mathbb{R}_{+}^{2}$ of the monomials
@@
The purpose of the latter construction is the following: $\partial C$ has to look like the graphs of different $P_{i}$’s at increasingly smaller neighbourhoods of the origin. Ultimately, this is achieved by ensuring that each $k_{i+1}$ is much larger than $k_{i}$. Indeed, for every $i\in\mathbb{N}$, mark by $(x_{i},y_{i})$ the left endpoint of $P_{i}$ after being assembled into $P$. Then, since all the curvatures $\kappa_{i}$ in (3.2) are uniformly bounded from above by $2x^{-1}$ for $x\in(0,1]$, it follows that we may get any $(x_{i},y_{i})$ arbitrarily close to the origin. The fact that $k_{i}$ may be chosen arbitrarily large ensures that we may get $P_{i+1}^{\prime}(b_{i+1})$ arbitrarily close to 0, and this, in turn, ensures that we may get the inner normal at $(x_{i},y_{i})$ arbitrarily close to $(0,1)$.
Then, we are interested in how the chords $K_{C}(\theta+\pi/2,\lambda)$ behave for small values of $\theta$. To this purpose, let $\beta\in(1,2)$, and consider a second planar convex body $A(\beta)$ whose boundary coincides with the graph of $y=|x|^{\beta}$ in $1,1]^{2}$. Moreover, suppose that $A(\beta)$ has a centre of symmetry and is symmetric with respect to the $y$-axis. Hence, it holds the following estimates on the chords of $A(\beta)$.
@@
We leave the proof of the latter in the Appendix, and we mention that similar calculations had previously been carried over in [beretti2025fouriertransformplanarconvex, Prop. 8.4].
Now, recall that, by a suitable choice of $k_{i}$, we may get $(x_{i},y_{i})$ arbitrarily close to the origin, as we may get its inner normal to make an angle arbitrarily close to $\pi/2$. Ultimately, this guarantees that, for every $i\in\mathbb{N}$, we may obtain an interval $\lambda_{i,1},\lambda_{i,2}]$ and a positive value $\theta_{i}$ such that, for every $\theta$ such that $|\theta|\leq\theta_{i}$ and every $\lambda$ such that $\lambda\in\left[\lambda_{i,1},\lambda_{i,2}\right]$, it holds
@@
Notice that, as $\theta_{i}$ and $\lambda_{i,2}$ may depend on $\left\{\beta_{j}\right\}_{j=1}^{i-1}$ and $\left\{k_{j}\right\}_{j=1}^{i-1}$, the choice of $\lambda_{i,1}$ can be made arbitrarily small at every $i$-th step; this will be key for the estimates in the following subsection.
Finally, Proposition 3.3 and Lemma 3.2 may be applied to $C$ to obtain the following result. From now on, we write $\rho^{-1}$ in place of $\lambda$, to prioritise the argument of the Fourier transform.
@@
In particular, once $\rho_{i,1}$ and $\theta_{i}$ are fixed, the endpoint $\rho_{i,2}$ can be made arbitrarily large.
@@ 3.2 Lower bound
To this purpose, we need an argument of Cassels [MR0087709] and Montgomery [MR1297543, Ch. 6] for estimating exponential sums from below. We formulate it in a version specialised to balls, and the reader may find a short proof either in [MR4358540] or [beretti2025fouriertransformplanarconvex]. In what follows, $B(r)$ stands for the ball centred at the origin with radius $r$.
@@
For every origin-symmetric convex body $C\subset\mathbb{R}^{2}$ and for every finite set of points $\{\mathbf{p}_{j}\}_{j=1}^{N}\subset\mathbb{T}^{2}$, it holds
@@
Now, in order to apply Cassels-Montgomery lemma effectively, we need to find a positive parameter $Z=Z(N)$ such that in the annulus $|\mathbf{m}|\in\left[\rho_{i,1},\rho_{i,2}\right]$ it holds
@@
Last, we consider the remaining region defined by $|\mathbf{m}|\in[Y,X]$ and $\|\arg(\mathbf{m})+\pi/2\|_{\pi}>1/10$. By some elementary geometry, we find that $\Phi(\mathbf{m})\leq 2\alpha$ where $\alpha$ is such that $|\mathbf{m}|\sin\alpha=Y/2$. Hence, it holds $\Phi(\mathbf{m})\leq\pi Y/(2|\mathbf{m}|)$, and since in this region we are aiming for $Z\Phi(\mathbf{m})\leq|\mathbf{m}|^{-3}$, then we settle for
@@
Starting from (3.1), and by Parseval’s identity and Cassels-Montgomery lemma, we get that, for any set of $N$ points $\mathcal{P}=\{\mathbf{p}_{j}\}_{j=1}^{N}\subset\mathbb{T}^{2}$, it holds
@@
Notice that, since $N^{-\varepsilon_{i}}$ can be made arbitrarily small, we can ultimately get the implicit constant before $N^{\alpha_{i}-\varepsilon_{i}}$ to be $1$.
@@ 3.3 Upper bound
This is done by employing suitable samplings, and we will make use of the kind that appear in [MR4358540, Thm. 8].
Fix $i\in\mathbb{N}$, and again, we remind that at every $i$-th step of our construction, $\rho_{i,2}$ serves as a free parameter that may be chosen arbitrarily large. By pairing Lemma 3.2 with (3.4), and by Lemma 3.4 and (3.3), in the annulus $\rho\in[\rho_{i,1},\rho_{i,2}]$ it holds the upper bound
@@
Since we have constructed $\partial C$ in such a way that its curvature is uniformly bounded from below, then by Lemma 3.2, for every $\rho>\lambda_{C}^{-1}$ and any angle $\theta$, it holds
@@
Set $N=GL$ (trivially, $N/2<\tilde{N}<2N$), and consider the set of $N$ points $\mathcal{P}\subset\mathbb{T}^{2}$ defined by
@@
For $j\in\{1,2,3\}$, notice that by the constraints we have imposed before, it holds $V_{j}\cap\mathcal{A}=V_{j}\cap B(\rho_{i,2})$ and $G^{2}L^{2}\mathcal{Q}_{i}=N^{2}\mathcal{Q}_{i}<1$, and by further applying (3.6), we obtain
@@
$\lesssim G^{-3}\sum_{n_{1}=1}^{+\infty}n_{1}^{-2}G/(\theta_{i}L)\lesssim G^{-2}L^{-1}/\theta_{i}.$
@@
$\mathcal{D}_{2}(\mathcal{P},\,C)\lesssim_{\beta_{i}}G^{2}L^{2}\left(L^{-2-2/\beta_{i}}+G^{-2}L^{-1}/\theta_{i}\right)\lesssim G^{2}L^{-2/\beta_{i}}+L/\theta_{i}\lesssim N^{\alpha_{i}}/\theta_{i}$
@@
whereas improper sums are conventionally considered as zeros. Now, notice that the latter definition implies that
@@
$N-\sum_{j=1}^{5}\lfloor n_{j}^{1-\alpha_{i}}\rfloor\,\lfloor n_{j}^{\alpha_{i}}\rfloor\leq 2^{10}N^{\left(1-\alpha_{i}\right)^{5}}\ll\bigl(N^{\alpha_{i}/2}\bigr).$
Then, if we further impose that $N^{\alpha_{i}^{2}/2}>\rho_{i,1}$, we may associate a choice of points as in the previous construction to every $N_{j}=\lfloor n_{j}^{1-\alpha_{i}}\rfloor\,\lfloor n_{j}^{\alpha_{i}}\rfloor$ such that $N_{j}\gtrsim N^{\alpha_{i}/2}$, and we don’t bother the contribution of the remaining points since, by the latter equation, it is certainly much smaller than $N^{\alpha_{i}}$. Finally, it is harmless to sum the contribution of our (at most) five point-sets and the remainder points, since, for $a_{1},\ldots,a_{6}\geq 0$, it holds
@@ Appendix A Auxiliary results
Before proceeding with the proof of Proposition 3.3, we first state the following auxiliary result. We will not present its proof here, as it is a simple exercise; however, the reader may find a proof in [beretti2025fouriertransformplanarconvex, Lemma 8.1].
@@
For the sake of brevity, we write $x_{0}=x_{o}(\theta+\pi/2)$ and $x_{+}=x_{+}(\theta+\pi/2,\lambda)$. Now, we notice that $x_{+}$ is the abscissa of the (right) intersection of the curve $y=|x|^{\beta}$ with the straight line
@@
Equalizing, and with the normalisation $z=\frac{x-x_{o}}{x_{o}}$, we get to the equation
@@
The conclusion hence follows by a rearrangement of the terms in (A.6). ∎

### v2

@@ 3 The second method
The harmonic analysis arguments we are to employ in this section have been recently developed in [BT22] and [BER25]. These make use of a geometric quantity that we define as follows.
@@
Let $C\subset\mathbb{R}^{2}$ be a convex body. For an angle $\theta\in[0,2\pi)$ and a value $\lambda>0$, we define the chord of $C$ in direction $\mathbf{u}(\theta)=(\cos\theta,\sin\theta)$ at distance $\lambda$ as
@@
One may notice that the decay of $|K_{C}(\theta,\lambda)|$ as $\lambda\to 0^{+}$ depends strongly on the local geometry of the (if just one) point $b\in\partial C$ whose inner normal is $\mathbf{u}(\theta)$. For example, if $\partial C$ is $\mathcal{C}^{2}$ and has positive curvature at $b$, then $|K_{C}(\theta,\lambda)|$ ultimately decays as $\lambda^{1/2}$, while if $b$ is the vertex at a corner, then $|K_{C}(\theta,\lambda)|$ ultimately decays as $\lambda$.
@@
We refer to [BER25, Lem. 3.3] for a detailed proof of the latter, but we mention that it relies upon the one-dimensional results in [POD91] and [BT22].
@@
${\mu}=\sum_{\mathbf{p}\in\mathcal{P}}\mu_{\rm D}(-\mathbf{p})-N\mu_{\rm H},$
@@
$\mathcal{D}(\mathcal{P},\,\boldsymbol{\tau}+C)=\int_{\mathbb{T}^{2}}\mathfrak{P}\{\mathds{1}_{C}\}(-{\boldsymbol{\tau}}-\mathbf{x})\,\mathrm{d}{\mu}(\mathbf{x})=\left(\mathfrak{P}\{\mathds{1}_{C}\}\ast{\mu}\right)(-\boldsymbol{\tau}).$
@@
the function of the Fourier coefficients of $f$. In particular, we note that it holds $\mathcal{F}\{{\mu}\}(\mathbf{0})=0$, and moreover, for every $\mathbf{n}\in\mathbb{Z}^{2}$, it holds
@@ 3.1 Geometric construction and Fourier estimates
We start our construction. Consider a planar convex body $C$ such that it has a centre of symmetry (this ensures $\gamma_{C}=|K_{C}|$) and is symmetric with respect to the $y$-axis. Then, suppose that $\mathbf{0}\in\partial C$ and that its inner normal is the vector $(0,1)$. By symmetry, we may construct $\partial C$ in the first quadrant $\mathbb{R}_{+}^{2}$. For this purpose, consider the graphs in $\mathbb{R}_{+}^{2}$ of the monomials
@@
The purpose of the latter construction is the following: $\partial C$ has to look like the graphs of different $P_{i}$’s at increasingly smaller neighbourhoods of the origin. Ultimately, this is achieved by ensuring that the sequence $\left\{k_{i}\right\}_{i\in\mathbb{N}}$ increases very rapidly. Indeed, for every $i\in\mathbb{N}$, mark by $(x_{i},y_{i})$ the left endpoint of $P_{i}$ after being assembled into $P$, as in Figure 3. First, the fact that, for $i\in\mathbb{N}$ and $x\in(0,1]$, it holds
$\kappa_{i}(x)\leq 2x^{-1}\quad\text{and}\quad P^{\prime}_{i}(x)\leq 2,$
ensures that we may get any $(x_{i},y_{i})$ arbitrarily close to the origin by a large enough choice of $k_{i}$. Secondly, since the elements in $\left\{k_{j}\right\}_{j\geq i}$ may be chosen arbitrarily large, we may get the elements in $\left\{P_{j+1}^{\prime}(b_{j+1})\right\}_{j\geq i}$ arbitrarily close to 0, and this, in turn, ensures that we may get the inner normal at $(x_{i},y_{i})$ arbitrarily close to the vector $(0,1)$; in particular, this prevents the curve segment $P$ from becoming a spiral.
In order to apply Lemma 3.2 to $C$, we are interested in how the chords $K_{C}(\theta+\pi/2,\lambda)$ behave for small values of $\theta$. For this purpose, let $\beta\in(1,2)$, and consider a second planar convex body $A(\beta)$ whose boundary coincides with the graph of $y=|x|^{\beta}$ in $1,1]^{2}$. Moreover, suppose that $A(\beta)$ has a centre of symmetry and is symmetric with respect to the $y$-axis. Hence, it holds the following estimates on the chords of $A(\beta)$.
@@
We leave the proof of the latter in the Appendix, and we mention that similar calculations had previously been carried out in [BER25, Prop. 8.4].
Recall that, by a large enough choice of the elements in $\left\{k_{j}\right\}_{j\geq i}$, we may get $(x_{i},y_{i})$ arbitrarily close to the origin, as we may get the inner normal at $(x_{i},y_{i})$ to be arbitrarily close to the vector $(0,1)$. Ultimately, this guarantees that, for every $i\in\mathbb{N}$, we may obtain an interval $\lambda_{i,1},\lambda_{i,2}]$ and a positive value $\theta_{i}$ such that, for every $\theta$ such that $|\theta|\leq\theta_{i}$ and every $\lambda$ such that $\lambda\in\left[\lambda_{i,1},\lambda_{i,2}\right]$, it holds
@@
Notice that, as $\theta_{i}$ and $\lambda_{i,2}$ depend on $\left\{\beta_{j}\right\}_{j=1}^{i-1}$ and $\left\{k_{j}\right\}_{j=1}^{i-1}$, the choice of $\lambda_{i,1}$ can be made arbitrarily small at every $i$-th step; this will be key for the estimates in the following subsection.
Finally, Proposition 3.3 and Lemma 3.2 may be applied to $C$ to infer the following result. From now on, we write $\rho^{-1}$ in place of $\lambda$, to prioritise the argument of the Fourier transform.
@@
In particular, although $\rho_{i,1}$ and $\theta_{i}$ are determined by the elements in $\left\{k_{j}\right\}_{j=1}^{i-1}$, the endpoint $\rho_{i,2}$ can be made arbitrarily large upon a larger choice of $k_{i}$.
@@ 3.2 Lower bound
For this purpose, we need an argument of Cassels [CAS56] and Montgomery [MON94, Ch. 6] for estimating exponential sums from below. We formulate it in a version specialised to balls, and the reader may find a short proof either in [BT22] or [BER25]. In what follows, $B(r)$ stands for the ball centred at the origin with radius $r$.
@@
For every origin-symmetric convex body $C\subset\mathbb{R}^{2}$ and for every finite configuration of points $\{\mathbf{p}_{j}\}_{j=1}^{N}$ in $\mathbb{T}^{2}$, it holds
@@
Now, in order to apply the Cassels-Montgomery lemma effectively, we need to find a positive parameter $Z=Z(N)$ such that in the annulus $|\mathbf{m}|\in\left[\rho_{i,1},\rho_{i,2}\right]$ it holds
@@
Lastly, we consider the remaining region defined by $|\mathbf{m}|\in[Y,X]$ and $\|\arg(\mathbf{m})+\pi/2\|_{\pi}>1/10$. By some elementary geometry, we find that $\Phi(\mathbf{m})\leq 2\psi$ where $\psi$ is such that $|\mathbf{m}|\sin\psi=Y/2$. Hence, it holds $\Phi(\mathbf{m})\leq\pi Y/(2|\mathbf{m}|)$, and since in this region we are aiming for $Z\Phi(\mathbf{m})\leq|\mathbf{m}|^{-3}$, then we settle for
@@
Starting from (3.1), and by Parseval’s identity and the Cassels-Montgomery lemma, we get that, for any configuration of $N$ points $\mathcal{P}=\{\mathbf{p}_{j}\}_{j=1}^{N}$ in $\mathbb{T}^{2}$, it holds
@@
Notice that, since $N^{-\varepsilon_{i}}$ can be made arbitrarily small, we can ultimately get the implicit constant in front of $N^{\alpha_{i}-\varepsilon_{i}}$ to be $1$.
@@ 3.3 Upper bound
This is done by employing suitable sampling sequences, and we will make use of the kind that appear in [BT22, Thm. 8].
Fix $i\in\mathbb{N}$, and again, we recall that at every $i$-th step of our construction, $\rho_{i,2}$ serves as a free parameter that may be chosen arbitrarily large. By pairing Lemma 3.2 with (3.4), and by Lemma 3.4 and (3.3), in the annulus $\rho\in[\rho_{i,1},\rho_{i,2}]$ it holds the upper bound
@@
By Lemma 3.2, and since we have constructed $\partial C$ in such a way that its curvature is uniformly bounded from below, then, for every $\rho>\lambda_{C}^{-1}$ and any angle $\theta$, it holds
@@
Set $N=GL$ (trivially, $L<G$ and $N/2<\tilde{N}<2N$), and consider the configuration of $N$ points $\mathcal{P}$ in $\mathbb{T}^{2}$ defined by
@@
For $j\in\{1,2,3\}$, notice that by the constraints we have imposed earlier, it holds $V_{j}\cap\mathcal{A}=V_{j}\cap B(\rho_{i,2})$ and $G^{2}L^{2}\mathcal{Q}_{i}=N^{2}\mathcal{Q}_{i}<1$, and by further applying (3.6), we obtain
@@
$\lesssim G^{-3}\sum_{n_{1}=1}^{+\infty}n_{1}^{-2}G/(\theta_{i}L)\lesssim G^{-2}L^{-1}\theta_{i}^{-1}.$
@@
$\mathcal{D}_{2}(\mathcal{P},\,C)\lesssim_{\beta_{i}}G^{2}L^{2}\left(L^{-2-2/\beta_{i}}+G^{-2}L^{-1}/\theta_{i}\right)\lesssim G^{2}L^{-2/\beta_{i}}+L/\theta_{i}\lesssim N^{\alpha_{i}}\theta_{i}^{-1}$
@@
where improper sums are conventionally considered as zeros. Now, notice that the latter definition implies that
@@
$N-\sum_{k=1}^{5}\lfloor n_{k}^{1-\alpha_{i}}\rfloor\,\lfloor n_{k}^{\alpha_{i}}\rfloor\leq 2^{10}N^{\left(1-\alpha_{i}\right)^{5}}\ll\bigl(N^{\alpha_{i}/2}\bigr).$
Then, if we further impose that $N^{\alpha_{i}^{2}/2}>\rho_{i,1}$, we may associate a choice of points as in the previous construction to every $N_{j}=\lfloor n_{j}^{1-\alpha_{i}}\rfloor\,\lfloor n_{j}^{\alpha_{i}}\rfloor$ such that $N_{j}\gtrsim N^{\alpha_{i}/2}$, and we do not bother the contribution of the remaining points since, by the latter equation, it is certainly much smaller than $N^{\alpha_{i}}$. In particular, it is not restrictive to admit repetition of points in the sampling sequence $\mathcal{P}_{N}$. Finally, it is harmless to sum the contribution of our (at most) five point-configurations and the remainder points, since, for $a_{1},\ldots,a_{6}\geq 0$, it holds
@@ Appendix A Auxiliary results
Before proceeding with the proof of Proposition 3.3, we first state the following auxiliary result. We will not present its proof here, as it is a simple exercise; however, the reader may find a proof in [BER25, Lemma 8.1].
@@
For the sake of brevity, we will write $x_{o}=x_{o}(\theta+\pi/2)$ and $x_{+}=x_{+}(\theta+\pi/2,\lambda)$. Now, notice that $x_{+}$ is the abscissa of the (right) intersection of the curve $y=|x|^{\beta}$ with the straight line
@@
Equating, and with the normalisation $z=\frac{x-x_{o}}{x_{o}}$, we get to the equation
@@
The conclusion follows by a rearrangement of the terms in (A.6). ∎